Concord Company uses special strapping equipment in its packaging business. The equipment was purchased in January 2019 for $10,500,000 and had an estimated useful life of 8 years with no salvage value. At December 31, 2020, new technology was introduced that would accelerate the obsolescence of Concord’s equipment. Concord’s controller estimates that expected future net cash flows on the equipment will be $6,615,000 and that the fair value of the equipment is $5,880,000. Concord intends to continue using the equipment, but it is estimated that the remaining useful life is 4 years. Concord uses straight-line depreciation.

A) Prepare the journal entry (if any) to record the impairment at December 31, 2020.

B) Prepare the journal entry for the equipment at December 31, 2021. The fair value of the equipment at December 31, 2021, is estimated to be $6,195,000.

C) Prepare the journal entry (if any) to record the impairment at December 31, 2020 and for the equipment at December 31, 2021, assuming that Concord intends to dispose of the equipment and that it has not been disposed of as of December 31, 2021.
Answer #1

a.

Carrying value of asset = 10,500,000 – [(10500,000/8) X2] = 10,500,000 – 2625,000 = $ 7,875,000

The depreciation for 2 years of $ 2625,000 has been subtracted from above.

Impairment loss = carrying value -fair value = 7875,000 – 5880,000 =1,995,000

Date

Particulars

Debit ($)

Credit ($)

Loss on impairment

1,995,000

         Accumulated depreciation

1,995,000

(To record impairment)

b.

This year, fair value has been revised. So, the depreciation needs to be adjusted as

Date

Particulars

Debit ($)

Credit ($)

Depreciation expense (6195,000/4)

1,548,750

         Accumulated depreciation

1,548,750

(To record depreciation)

c.

Date

Particulars

Debit ($)

Credit ($)

For a

Loss on impairment

1,995,000

         Accumulated depreciation

1,995,000

(To record impairment)

For b

Accumulated depreciation

315,000

          Recovery of impairment loss (6195,000 – 5880,000)

315,000

(To record recovery of impairment loss)

kindly upvote